OpenStudy (thomasfire55):

Solve for x. (2x - 1)(3x + 2) = x + 292

OpenStudy (robtobey2):

Expand the LHS and then move the RHS to the LHS and simplify. There is still more to do.

OpenStudy (robtobey2):

Expand means to multiply out.

OpenStudy (thomasfire55):

?

OpenStudy (princeharryyy):

are you given options?

OpenStudy (thomasfire55):

no

OpenStudy (robtobey2):

The following is from the Mathematica program and the terms are not in classroom order. \[6 x^2+x-(x+292)-2=0 \]\[6 x^2-294=0\]The above can be simplified and factored.

OpenStudy (princeharryyy):

than you have to multiply the LHS and solve for the value of x.

jhonyy9 (jhonyy9):

(2x-1)(3x-2) =x+292 6x^2 -4x-3x +2 =x+292 6x^2 -8x -290 =0 3x^2 -4x -145 =0 so just solve this quadratic for x

OpenStudy (radar):

@jhonyy9 You did not copy the problem as given, you changed the sign.

OpenStudy (robtobey2):

\[6 x^2-294=0 \]\[x^2-49=0 \]

jhonyy9 (jhonyy9):

yes right there is a sign inversed what i changed - sorry

OpenStudy (radar):

yes, robtobey has resolved it down to the difference of two perfect squares, where it is easy to solve by factoring.